OpenStudy (anonymous):

(x+9)(x-9) is this the same as the FOIL method, I'm lost

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

the answer is x^2-81

OpenStudy (anonymous):

first = x^2 outer = -9x inner = +9x last = -81 the outer and inner will be canceld out so the answer is x^2-81 :-)

OpenStudy (anonymous):

That can be recognized as "the difference of squares".. There is a formula..

OpenStudy (anonymous):

And you can "foil" it.

OpenStudy (anonymous):

(x-a)(x+a) = x^2 - a^2 if you foil it, it will turn out the same because two of the terms end up to cancel...